OpenStudy (anonymous):

use the derivative of cos^2(x) to find the integral:

OpenStudy (anonymous):

and where's the integral ?

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \sin2x }{ \cos^2(x) + 9 }\]

OpenStudy (anonymous):

WHOOPS forgot the dx

OpenStudy (anonymous):

aaaaaand, i think i've got the ans. But if anyone else has it, i'd like to check anyways :P

OpenStudy (anonymous):

sin2x = 2sinx cosx cos x =t -sin x dx =dt then substitute t^2 as v You get 2t as dv

OpenStudy (anonymous):

nono we don't do that, we put a minus sign inside and outside the integral sign, and then we substitute -sin2x dx for dy, if we say that cos^2(x) = y

OpenStudy (anonymous):

and then the denominator is y + 9 = y +3^2, so it becomes cot'(y) + c

OpenStudy (anonymous):

it's the same thing however

OpenStudy (anonymous):

you'll get the answer in log not in cot x

OpenStudy (anonymous):

no we wont, the derivative of cotINVERSEx is -1/1+x^2 so it'll be that into 1/3

OpenStudy (anonymous):

you'll get -2t/t^2 + 9 as your integral after substituting cos x = t and t/t^2 + 9 is not cot inverse x

OpenStudy (anonymous):

it's not t, it's dt.

OpenStudy (anonymous):

ughh how do i explain this poor fellow well dude, why do you ask questions here when you know you'll get the correct answer ? i've checked it twice and thrice and what i'm saying is the answer

OpenStudy (anonymous):

ok, ok, ok, i got it, sorry. My bad.

OpenStudy (anonymous):

fine